How many ions are present in a 25mL solution of 30mM CaSO4?
Answer #1

CaSO4 Undegoes dissociation to give rise to Ca+2 and SO4-2. Hence there will be 2 ions per molecule ( Ca +2 ion and SO4-2 ion)

30mM CaSO4 =30*10-3 M CaSO4 =

Volume =25 ml , hence 25 ml corresponds to (25/1000) *30X10-3 = 7.5X10-4 moles

number of ions per moles =2

total number of ions =7.5X10-4*2 =1.5X10-3 ions/moles

1 mole of any substance cosist of 6.023X1023 molecules, hennce total number of ions =1.5X10-3 * 6.023*1023 = 9.0345*1020